OpenStudy (anonymous):

Multiply and simplify : ( 5 + 2v2) ( 8 - 3v2 ) V stands for sqrt A. 40 - 5v2 B. 28 + v2 C. 46 + 2V2 I got A.

OpenStudy (anonymous):

\[(5+2\sqrt{2})(8-3\sqrt{2})=40-15\sqrt{2}+16\sqrt{2}-6\sqrt{2}*\sqrt{2}=40+\sqrt{2}-12=28+\sqrt{2}\] FOIL

OpenStudy (anonymous):

\[28+\sqrt{2}\] may have been cut off